OpenStudy (anonymous):

Simplify the following expression. a + 3c − 4b − c + a − 2b

OpenStudy (anonymous):

add like terms to simpify!

OpenStudy (anonymous):

u will earn a metal!!!!

OpenStudy (anonymous):

a+a=? 3c+c=? -4b+2b=? then you'll have it! :)

OpenStudy (anonymous):

if u answers the question

OpenStudy (anonymous):

I don't get out answers, I help those get the answers! sorry

OpenStudy (anonymous):

a+3c-4b-c+a-2b 2a+2c-6b

OpenStudy (anonymous):

tin tin tin correct

OpenStudy (anonymous):

ok well you start out by pairing the variables so it would be a+a and 3c -c and -4b-2b the answer would be 2a-6b-2c

OpenStudy (anonymous):

2a − 6b + 2c

OpenStudy (anonymous):

ya thats what i ment

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

u earn a metal

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

lol your welcome